OpenStudy (anonymous):

Which of the following represents a linear function ? A. f(x)=3^x B. f(x)=x^3 C. f(x)=3x^2 D. f(x)=x+3

OpenStudy (anonymous):

the answer is B. f(x)=x^3

OpenStudy (freckles):

x^3 is not linear

OpenStudy (freckles):

f(x)=mx+b is a linear function

OpenStudy (freckles):

f(x)=a^x is exponential for 0<a<1 union a>1 f(x)=ax^3+bx^2+cx+d is cubic polynomial f(x)=ax^2+bx+c is a parabola so what do you think the answer is ?

OpenStudy (anonymous):

the answer is d because x is a constant growth or the slope, and 3 is the initial start

OpenStudy (anonymous):

thank you all ;)